martha walked to the park which is 12 blocks from her home and jogged back. the entire trip took 1.5 hours. if martha jogged three times faster than she walked how fast did she walk?
Answer by oberobic(2304)

The whole trip took 1.5 hrs.
.
d=rt, so t = d/r
.
1.5 = 12/x + 12/3x
.
Multiply both sides by 3x.
.
3x(1.5) = (3x*12)/x + (3x*12)/3x
4.5x = 36 + 12
4.5x = 48
x = 10.667 or 10+2/3
.
So, this suggests her walking speed is 10+2/3 blocks/hr and her jogging speed is 32 blocks/hr.
.
Checking our work,
12 = 10.667t
10.667t = 12
t = 12/10.667
t = 1.125 hr walking
.
12 = 32t
32t = 12
t = 12/32 = .375 hr jogging
.
1.125 + .375 = 1.5 hr total
.
OK